You are given the following three points A(1,−3,5), B(4,-11,1) and C(-3,8,5).

Which of the following statements are true?


  1. The vector equation of the line AB is "\mathbf{r}=(1,-3,5)+\lambda (3,-8,-4)" , where "\lambda" is a parameter.
  2. The three points  A, B  and C ie on the same straight line
  3. "\underline{AB} \times \underline{AC}=(44,16,1)"
  4. Lines AB and AC are perpendicular.
  5. The equation of a plane passing through the points AB and C is "44x+16y+z=1"
1
Expert's answer
2021-10-18T07:40:42-0400

 Given three points are  A(1,−3,5), B(4,-11,1) and C(-3,8,5).

Let us check every statement true or false individually.

Statement-1.

Vector equation of line AB is

"\textbf{r}" = (1,-3,5) + "\lambda" (4-1, -11+3,1-5)

=> "\textbf{r}" = (1,-3,5) + "\lambda" (3, -8,-4)

So Statement-1 is "\textbf{true}" .

Statement-2.

The three points A, B and C will lie on the same straight line if position vector of C satisfy the equation of AB.

i.e. if (-3,8,5)=(1,-3,5)+"\lambda" (3,-8,4)

i.e if (-3,8,5)=(1+3"\lambda" ,-3-8"\lambda" ,5+4"\lambda" )

i.e if -3 = 1+3"\lambda" , 8=-3-8"\lambda" ,5=5+4"\lambda"

i.e if "\lambda" = "-\frac{4}{3}" , "\lambda" ="-\frac{11}{8}" , "\lambda" =0

As values of "\lambda" are different, position vector of C doesn't satisfy the equation of line AB.

So statement-2 is "\textbf{false}"

Statement-3.

"\overrightarrow{AB}" = (4-1, -11+3,1-5)=(3,-8,-4)

"\overrightarrow{AC}" = (-3-1,8+3,5-5)=(-4,11,0)

"\overrightarrow{AB}" X"\overrightarrow{AC}" = "\begin{vmatrix}
 \hat{i} &\hat{j} &\hat{k}\\
 3 & -8&-4\\-4&11&0
\end{vmatrix}"

= "\hat{i}" (44)+"\hat{j}" (16-0)+"\hat{k}" (33-32) = (44, 16, 1)

So Statement-3 is "\textbf{true}" .

Statement-4.

If line AB and line AC are perpendicular then "\overrightarrow{AB}." "\overrightarrow{AC}" = 0.

"\overrightarrow{AB}" ."\overrightarrow{AC}" = (3,-8,-4).(-4,11,0)

=-12-88-0=-100 ≠0

So line AB is not perpendicular to line AC

So Statement-4 is "\textbf{false}"

Statement-5

As every linear equation of x, y, z represents a plane and one and only one plane can be drawn through three non collinear points, equation of plane through A,B and C will be 44x+16y+z=1

If co-ordinates of the points A, B,C satisfy the equation of plane 44x+16y+z=1 .

For point A, 44.(1)+16.(-3)+1.(5)=1

=> 44-48+5=1

=> -4+5=1

=> 1=1

So A lies on the given plane.

For point B, 44.(4)+16.(-11)+1.(1)=1

=> 176-176+1=1

=> 1=1

So B lies on the given plane

For point C, 44.(-3)+16.(8)+1.(5)=1

=> -132+128+5=1

=> 1=1

So C lies on the given plane

So equation of plane through A, B, C is 44x+16y+z=1 .

Therefore Statement-5 is "\textbf{true}" .